# DiPEx: Dispersing Prompt Expansion for Class-Agnostic Object Detection

### v1

@@ 1 Introduction
To better understand the limitation of modern VLMs in generic object localization, we investigated the design of hand-crafted text queries (Section 2) to enhance detection recall through two approaches: (1) We employed a Universal query, using ChatGPT to generate 13 types of broad nouns and adjectives (e.g., "objects," "generic") as queries for the Grounding DINO model, aiming to detect a wide array of objects without focusing on specific categories; (2) We implemented a Class-wide query, selecting 25 high-level semantic words (e.g., "plant," "animal") from the top layer of the WordNet hierarchy (also used for the ImageNet vocabulary) to cover extensive object categories. Our findings, depicted in Figure 1b and Table 1, reveal that while VLMs can generalize across universal object categories, combining all queries into one string significantly reduces detection performance (by up to 52% in AR) due to the “semantic overlap” among words. This suggests that optimal detection requires conducting multiple separate inferences, presenting substantial computational demands for large datasets.
@@
To overcome the aforementioned limitations, we propose a novel self-supervised Dispersing Prompt Expansion (DiPEx) strategy. This approach progressively expands a set of non-overlapping hyperspherical prompts for capturing all objects in a given dataset, thereby benefiting downstream tasks such as out-of-distribution object detection. Specifically, we start with a generic parent prompt that is self-supervised using the Universal and Class-wide text queries. To capture more fine-grained semantics, we split the parent prompts with high semantic uncertainty into a set of distinct child prompts. We initialize child prompts by diversifying the parent token embedding, randomly rotating it to different angles on the hypersphere to yield a range of unique prompts. Dispersion losses are employed to minimize semantic overlap among child prompts while maintaining semantic consistency across parent-child prompt pairs. To prevent excessive growth of the prompt sets, we estimate the maximum angular coverage (MAC) of the semantic space as a criterion to terminate the prompt expansion process, balancing semantic richness and computational overhead. Extensive experiments on the MS-COCO and LVIS datasets verify the effectiveness and versatility of the proposed DiPEx strategy. With a single pass of inference, DiPEx can achieve by up to 20.1% improvements in average recall (particularly 35.2% for small objects) and outperforms segment anything model (SAM) [25] by 21.3% in average precision. The source code is available at https://github.com/jason-lim26/DiPEx.
Table 1: Zero-shot class-agnostic object detection performance of Grounding DINO [32] on MS-COCO [31], with hand-crafted prompts from various sources. We report average recall (AR) and precision (AP) limited to a maximum of 100 detections per image. $\Delta$AR quantifies the percentage decrease in AR comparing “query-merging” to “prediction-merging” for forming multi-world queries.
@@
Related Study. The full discussions can be found in Section A.1. Traditional bottom-up approaches for region proposal generation, such as those by [46] and [26], often face precision constraints despite high recall rates, limiting their scalability. Recent advancements in Vision Transformers (ViTs) by [4] and [9] have enabled self-supervised learning on massive datasets, extracting semantically meaningful features. Methods like LOST [44] and TokenCut [50] use graph-based techniques but are limited to detecting a single object per image. MOST [42] addresses this with entropy-based box analysis but struggles with generalization. MAVL [35] uses a late fusion strategy with text queries, requiring full supervision and multiple inferences. Our approach eliminates the need for labels and achieves state-of-the-art performance with one-pass inference using non-overlapping prompts. Vision-Language Models (VLMs), like those by [41] and [19], have shown potential in learning generic concepts. HierKD [34] and OV-DETR [56] align image representations with captions and extend DETR to open-vocabulary settings. GLIP [28], Grounding DINO [32], and T-Rex2 [20] integrate object detection and visual grounding. However, VLMs’ effectiveness depends on textual cues, and prompt tuning, as introduced by CoOp [23] and improved by CoCoOp and MaPLe [24], offers a solution by optimizing soft prompts while keeping the model’s parameters frozen. ProDA [33] learns diverse prompts using a Gaussian model. Despite these advancements, full supervision is typically required. UPL [16] and POUF [45] introduced unsupervised prompt learning, but adaptation for object detection remains limited. DiPEx is the first to apply prompt learning to class-agnostic object detection through a self-training approach.
2 Pilot Study: Hand-crafted Queries for Class-agnostic OD
In this section, we detail our preliminary exploration of the zero-shot detection capacities using one of the most state-of-the-art VLMs, Grounding DINO [32], to detect all objects irrespective of the associated classes on the MS-COCO dataset [31] as illustrated in Figure 1b. We consider the following two text queries:
@@ 2 Pilot Study: Hand-crafted Queries for Class-agnostic OD
Universal Query. We employ ChatGPT to generate 13 synonyms of universal concepts, including nouns and adjectives, which are displayed as x-axis labels. The zero-shot object detection results, measured by average recall (AR) and precision (AP) across the top 100 confident boxes for each query text, are presented. The plot reveals that more general terms such as "generic" and "items" yield the highest AR. Surprisingly, more specific descriptors like "foreground", "small", or "tiny" tend to reduce AR and do not effectively aid in identifying foreground or small objects.
@@
Discussion on Multi-Word Queries. The zero-shot results presented in Figure 1b are obtained using single-word prompts for the Grounding DINO. To explore whether combining multiple words as prompts from a given source (e.g., WordNet) could improve zero-shot detection performance, we developed strategies for merging at both the input stage (query-merging) and the output stage (prediction-merging) as shown in Table 1. The query-merging strategy concatenates all input text queries (e.g., "foreground. elements. $\cdots$ tiny. objects.") and performs a single-pass inference to obtain detections. The prediction-merging strategy, on the other hand, uses each text query individually for separate inference and then combines all box predictions. Table 1 shows that applying query-merging to Universal words results in a 52.46% reduction in AR compared to prediction-merging, whereas Class-wide queries (e.g., from WordNet) achieve a smaller decrease in AR of only 23.64%. Our findings suggest that large semantic overlaps in concatenated queries (e.g., "stuff", "objects" and "item" from ChatGPT) may greatly contribute to diminished OD performance. Another case study to support this observation is also presented in Section A.2. Consequently, we hypothesize that if we could develop a method to learn a set of semantically non-overlapping prompts for the target dataset, we could efficiently locate all objects with one-pass inference using VLMs.
@@ 3 Proposed Approach
In this section, we first mathematically formulate the task of class-agnostic detection using a general VLM and, without loss of generality, illustrate the process using Grounding DINO [32] as an exemplar model. We detail the steps of the proposed dispersing prompt expansion in Sec. 3.2, followed by the early termination strategy of the prompt set growth.
@@ 3.1 Problem Formulation
Class-agnostic OD. Let $\mathbf{I}$ denote the input image and $\mathbf{T}$ the associated text query. For the zero-shot object detection in a class-agnostic setting, we consider the text query $\mathbf{T}$ to be of the form of "a photo of a {class}", where the class token {class} is sampled from our predefined Universal (e.g., "objects") or Class-wide (e.g., "plant") sets as described in Section 2. The text query is then tokenized and projected into word embeddings as $\mathbf{P}=\{\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{M},\mathbf{c}\}$, where $\mathbf{v}=\{\mathbf{v}_{i}\}_{i=1}^{M}\in\mathbb{R}^{M\times d}$ indicates a set of $M$ contextual embeddings and $\mathbf{c}$ is the query text embedding. Here, $d$ indicates the dimensions of learnable tokens. The visual embeddings $\mathbf{E}_{v}$ extracted from the visual encoder and prompt embeddings $\mathbf{P}$ are fused jointly to prompt the VLM and generate the final bounding box predictions $\mathbf{O}=f(\mathbf{E}_{v},\mathbf{P})\in\mathbb{R}^{N_{B}\times 4}$, with $f$ being the VLM, and $N_{B}$ being the number of predicted boxes. Formally, the objective of class-agnostic OD is to ensure that the generated bounding boxes can capture any objects as comprehensively as possible.
@@
Figure 2: An illustration of the (1) proposed prompt expansion strategy that selectively grows a set of child prompts for the highlighted parent prompt across $L$ iterations; (2) diversifying initialized embeddings of the child prompt on a hypersphere and (3) quantifying maximum angular coverage $\alpha_{\operatorname{max}}$ for early termination of the prompt growth.
@@ 3.2 Dispersing Prompt Expansion (DiPEx)
4 Experiments
4.1 Experimental Setup
Datasets and Evaluation Metrics. We conduct our experiments using two detection datasets: (1) MS-COCO [31] is a large-scale object detection and instance segmentation dataset, containing about 115K and 5K images of 80 classes in the training and validation split, respectively. We test our model in a class-agnostic manner on the COCO val2017 split. (2) LVIS [14] collected 2.2 million high-quality instance segmentation masks with a broader spectrum of 1000 class labels, which naturally constitutes the long-tailed data distribution. It generally includes 100k training images and 19.8k validation images. To further validate the efficacy of DiPEx in downstream out-of-distribution object detection (OOD-OD) tasks, we evaluate our method using an rectified version of the OOD-OD benchmark. Unlike previous benchmarks [11], where samples that do not contain ID instances are manually selected and ID and OOD performance are evaluated separately, we tested our approach on the MS-COCO, which includes a mixture of both ID and OOD objects. While we followed the settings outlined in OOD-OD [11], with 20 base classes in VOC [12] designated as ID classes and the remaining classes treated as OOD, our choice of dataset provides a more rigorous evaluation by incorporating both ID and OOD instances together, providing a more rigorous assessment of our method’s performance in real-world scenarios.
Evaluation Metrics. We report class-agnostic object detection results on LVIS val split, containing about 5K images. For evaluation, we adopt official metrics from the COCO 2017 challenge. Specifically, we report average precision (AP) at IoU thresholds from 0.5 to 0.95 and average recall (AR) at the threshold range of [0.5:0.95] and across the scales - AR@S, AR@M, and AR@L for small, medium, and large objects, respectively. The implementation details of baselines and the proposed approach can be found in Appendix A.3.
Method Description AR1 AR10 AR100 AR@S AR@M AR@L AP
Selective Search [46] non-parametric 0.1 1.1 7.8 0.9 7.2 20.7 0.1
UP-DETR [7] self-training 0.0 0.0 0.4
DETReg [1] self-training 0.6 3.6 12.7
FreeSOLO [48] self-training 4.6 11.4 15.3
Exemplar-FreeSOLO [17] self-training 8.2 13.0 17.9
MOST [42] self-training 3.1 6.4 6.4 0.1 1.6 24.5 3.3
CutLER [49] self-training 6.8 19.6 32.8 13.7 37.5 60.0 29.6
Grounding DINO ["generic"] [8] zero-shot 10.3 37.8 44.1 17.7 51.6 80.0 28.3
DiPEx self-training 10.5 40.8 63.2 39.2 74.3 89.8 35.9
4.2 Main Results on Class-agnostic OD and OOD-OD
Class-agnostic OD on MS-COCO. To validate the proposed method for class-agnostic object detection, we compared it with ten different baseline methods on the MS-COCO dataset, using various metrics as reported in Table 2. We observed that non-parametric methods generally underperform compared to self-training methods due to their inability to learn and extract semantic and geometric information about objects from the dataset. In contrast, Grounding DINO, leveraging pre-trained knowledge, demonstrates strong zero-shot capabilities and achieves AR100 of 44.1% with a single text prompt “generic”. Furthermore, CoOp, which fine-tunes prompts for Grounding DINO, enhances class-agnostic detection performance by 3.4% in AR10 compared to direct zero-shot inference. Our method, which expands the learnable prompts to a wider angular distance, exceeds all baselines, achieving the highest performance across all metrics and outperforming the leading baseline, CoOp, by 4.3% in AR10. Notably, for small objects that are challenging to localize, our method improves AR@S by 7.7% compared to CoOp, indicating that expanded prompts better capture a ranger size of objects. Additionally, the proposed DiPEx achieved the highest AP of 35.9%, demonstrating the highest quality of class-agnostic detection.
Class-agnostic OD on LVIS. To further validate the efficacy of DiPEx, we conducted extensive experiments on a more challenging dataset, LVIS, which contains thousands of long-tail distributed class labels. As shown in Table 3, prompt tuning methods such as CoOp [23] and CoCoOp [22] yield better results than zero-shot Grounding DINO using hand-crafted prompts (e.g., "items," "generic," "objects"). CoCoOp also outperforms multi-object discovery baselines, such as CutLER [49] and HASSOD [2], by 86.7% and 51.3% in AR200, respectively. Notably, SAM [25], which was pre-trained on a vast amount of data (millions of images and billions of masks), shows strong zero-shot ability, surpassing all other baselines. Regarding the proposed DiPEx, it even outperforms SAM by 13.3% in AR200 and 21.3% in AP with a mere 4 epochs of self-training, and further outperforms CoOp by 20.1% in AR200.
Method AR1 AR10 AR200 AR@S AR@M AR@L AP@S AP@M AP@L AP
Selective Search [46] 0.1 1.1 13.0 6.1 19.9 37.6 0.2 0.4 0.2 0.2
G-DINO ["object"] [8] 4.1 17.9 27.2 13.0 44.1 71.1 5.6 10.0 9.4 5.4
G-DINO ["generic"] [8] 3.8 16.5 20.2 6.5 34.5 67.7 4.1 17.4 30.7 9.0
G-DINO ["items"] [8] 4.0 17.8 28.0 13.9 45.3 70.7 6.3 19.6 32.0 11.6
SAM [25] 42.7 27.7 66.3 75.5 6.1
CutLER [49] 2.4 9.3 21.8 10.8 35.1 55.5 2.7 9.1 15.1 4.5
HASSOD [2] 26.9 15.6 42.2 56.9
G-DINO + CoOp [23] 4.2 19.1 40.3 23.6 63.5 83.5 8.3 23.7 32.3 14.0
G-DINO + CoCoOp [22] 4.2 19.2 40.7 24.1 63.8 84.1 8.1 22.4 30.1 13.6
DiPEx 4.3 20.1 48.4 31.9 72.6 88.2 9.3 25.3 32.8 15.2
Method Known Unknown
AP AP50 AR100 AR@S AR@M AR@L AP AP@S AP@M AP@L
Selective Search [46] 8.3 1.0 8.5 23.2 0.1 0.0 0.0 0.5
MOST [42] 5.3 0.1 1.3 22.5 0.4 0.1 0.4 1.2
CutLER [49] 34.5 15.8 41.5 62.7 5.7 2.3 6.9 13.7
VOS [11] 36.6 56.7 10.0 2.2 6.1 27.1 2.8 0.8 2.2 7.2
PROB [60] 28.2 43.8 14.2 2.0 11.9 43.5
UnSniffer [29] 35.8 55.8 20.6 11.8 19.9 34.8 2.9 1.5 3.1 5.3
G-DINO ["generic"] 46.3 59.7 43.3 18.0 52.1 82.6 12.5 6.9 17.8 25.7
DiPEx 46.3 59.7 5.99 35.8 72.9 89.7 15.7 9.7 21.8 25.2
Downstream OOD-OD on MS-COCO. To verify the generalization capacity of the proposed DiPEx in the downstream task of out-of-distribution object detection (OOD-OD), we conducted experiments comparing the detection results of DiPEx for both known and unknown classes against extensive baselines. As shown in Table 4, the zero-shot Grounding DINO using known class names as prompts, combined with a simple “generic” prompt for unknown, significantly outperforms all other non-VLM methods (e.g., 25.5% higher AR100 compared to CutLER [49]). This is because VLMs benefit greatly from the rich semantic knowledge extracted by the language model, to better comprehend the object information in images. In terms of the proposed DiPEx, it learns to expand the text prompts in embedding space, enabling the capture and differentiating of objects with varying sizes and diverse semantics from known classes. This results in a significant performance improvement, achieving a 38.3% increase in AR100 and a 25.6% increase over zero-shot predictions. Furthermore, in scenarios with different known vocabularies, the expanded prompts can be directly applied to detect unknown objects, eliminating the need for retraining.
4.3 Ablation Study and Model Analysis
We investigate the impact of various factors on prompting performance including the learnable prompt lengths, the number of expansion rounds $L$, and angular coverage achieved across rounds. To facilitate model analysis, we present the distribution of prompt logit activation and visualization of detection results. Furthur ablation studies refers to Section A.4.
Impact on Number of Prompts. In Figure 3, we compare the impact of prompt length $N$ for DiPEx against CoOp [23] & CoCoOp [22]. Overall, we observe a consistent improvement in performance as the number of prompts increases, suggesting that a larger number of prompts allows for the capture of more comprehensive semantics. Interestingly, while CoOp’s [23] performance remains constant, CoCoOp’s [22] performance declines, suggesting that increasing the number of prompts does not necessarily guarantee enhanced performance.
Qualitative Study. In this section, we present visualized class-agnostic box predictions on images sampled from the MS-COCO dataset [31], as shown in Figure 6. The proposed DiPEx method demonstrates a superior ability to detect more bounding boxes than all baseline methods, particularly for small objects. For example, people in the distance (rows 1 and 3) and some bonsai (row 2) are missed by all baselines but successfully detected by DiPEx, showcasing its strong capability in localizing challenging small objects. For large objects, such as a motorcycle (row 3) and two people shaking hands in the near distance (row 1), DiPEx localizes them with significantly higher confidence compared to the zero-shot predictions of Grounding Dino using the prompt "generic". Additionally, DiPEx successfully identifies objects that are not annotated in the MS-COCO ground truth, such as plates (row 1), a pillowcase (row 2), and a frame on the wall (row 2). This highlights DiPEx’s ability to identify a comprehensive set of class-agnostic objects, even those missed in human annotations.
5 Conclusion and Limitations
Appendix A Appendix / Supplemental Material
This supplementary material includes a comprehensive overview of related work on class-agnostic object detection, vision-language models (VLMs), and prompt tuning. A case study illustrates the reduction in confidence when text queries semantically overlap. Additionally, detailed descriptions of baselines, implementation details for both the baselines and the proposed method, and an extensive ablation study are provided. The ablation study analyzes the impact of pseudo-labeled supervision and the effect of the hyperparameter $\gamma$. Lastly, we present a detailed algorithm and additional visualizations of class-agnostic box predictions.
  • Sec. A.1: Related Work
  • Sec. A.2: Case Study - Confidence Diminishing
  • Sec. A.3: Baselines and Implementation Details
  • Sec. A.4: More Ablation Study
  • Algorithm 1: Detailed Algorithm of DiPEx
  • Figure 9 and 10: Additional Visualizations of Class-Agnostic Box Predictions
A.1 Related Work
Class-Agnostic Object Detection. Traditional bottom-up approaches [46, 26, 59, 39, 40, 6] for region proposal generation, often grapple with the constraints precision, despite high recall rates, reducing their scalability for general use in diverse environments. Recent breakthroughs in ViTs [4, 9, 38] have enabled scaling up to massive datasets for self-supervised learning, extracting both local and global semantically meaningful features. This has led to numerous methods in unsupervised object discovery and localization. LOST [44] is an early application, using a patch similarity graph and an inverse degree map to identify seed patches and extract bounding boxes. TokenCut [50] constructs an undirected graph with image tokens as nodes, applying the normalized cut algorithm [43] for foreground-background segregation. MOVE [36] builds on LOST by employing deep spectral bipartitioning, offering a more principled and effective approach. However, both LOST [44] and TokenCut [50] are limited to detecting a single object per image. MOST [42] addresses this limitation by using entropy-based box analysis (EBA) to segregate foreground tokens. Nevertheless, their performance remains sub-optimal, constrained by their limited capacity to generalize across diverse object categories. Closest to our work is MAVL [35], where they develop an MViT with late fusion strategy and use generic text queries like "all objects" to locate objects. However, their framework requires full supervision and multiple inferences with different textual prompts, yet still falls short of achieving optimal performance. In contrast, our approach eliminates the need for labels and achieves SOTA performance with one-pass inference with the non-overlapping prompts.
A.2 Case Study: Confidence Diminishing when Text Query Semantically Overlap
Case Study. To verify our hypothesis made in Section 1, we conduct a case study to show the phenomenon of confidence diminishing with multi-word prompts. To quantify the semantic overlap and its influence on zero-shot OD performance, we calculate the angular distance between pairs of textual tokens generated by BERT [8]. As shown in Figure 7, a small angular distance $\theta$ of 53.73 between the text tokens “plates” and “dishes” diminishes the model’s confidence. Consequently, some boxes that could be precisely localized with high confidence using the single token “plates” are omitted. In contrast, concatenating two text tokens with a large angular distance (e.g., 60.99 between “plates” and “cup") results in confident box predictions that cover all the boxes predicted with each individual token (“plates” or “cup”).
A.3 Baselines and Implementation Details
Implementation Details. Our code is developed on the Open Gounding-DINO framework [61], and operates on a single NVIDIA RTX A6000 GPU with 48 GB of memory. For our experiments, we choose a batch size of 8 for training, and set hyperparameter $\gamma=0.1$, $\tau_{p}=0.01$, $\tau_{c}=0.01$, $\theta=5^{\circ}$, $K=9$, $L=3$, and while adopting all remaining hyperparameters from the Open Grounding-DINO codebase. We empirically set the $\mathcal{T}_{\alpha}=75^{\circ}$ as our threshold for expansion termination. The original implementation of CoOP was developed for image classification tasks based on CLIP and supervised contrastive learning. We extend CoOP to class-agnostic object detection using pseudo labeling-based self-training, which remains consistent with our approach. All the implementation code and configurations files are provided in supplementary materials and will be publicly released upon acceptance of this work.

### v2

@@ 1 Introduction
To better understand the limitation of modern VLMs in generic object localization, we investigated the design of hand-crafted text queries (Section 2) to enhance detection recall through two approaches: (1) We employed a Universal query, using ChatGPT to generate 13 types of broad nouns and adjectives (e.g., “objects”, “generic”) as queries for the Grounding DINO model, aiming to detect a wide array of objects without focusing on specific categories; (2) We implemented a Class-wide query, selecting 25 high-level semantic words (e.g., “plant”, “animal”) from the top layer of the WordNet hierarchy (also used for the ImageNet vocabulary) to cover extensive object categories. Our findings, depicted in Figure 1b and Table 1, reveal that while VLMs can generalize across universal object categories, combining all queries into one string significantly reduces detection performance (by up to 52% in AR) due to the “semantic overlap” among words. This suggests that optimal detection requires conducting multiple separate inferences, presenting substantial computational demands for large datasets.
@@
To overcome the aforementioned limitations, we propose a novel self-supervised Dispersing Prompt Expansion (DiPEx) strategy. This approach progressively expands a set of non-overlapping hyperspherical prompts for capturing all objects in a given dataset, thereby benefiting downstream tasks such as out-of-distribution object detection. Specifically, we start with a generic parent prompt that is self-supervised using the Universal and Class-wide text queries. To capture more fine-grained semantics, we split the parent prompts with high semantic uncertainty into a set of distinct child prompts. We initialize child prompts by diversifying the parent token embedding, randomly rotating it to different angles on the hypersphere to yield a range of unique prompts. Dispersion losses are employed to minimize semantic overlap among child prompts while maintaining semantic consistency across parent-child prompt pairs. To prevent excessive growth of the prompt sets, we estimate the maximum angular coverage (MAC) of the semantic space as a criterion to terminate the prompt expansion process, balancing semantic richness and computational overhead. Extensive experiments on the MS-COCO and LVIS datasets verify the effectiveness and versatility of the proposed DiPEx strategy. With a single pass of inference, DiPEx can achieve by up to 20.1% improvements in average recall (particularly 35.2% for small objects) and outperforms segment anything model (SAM) [26] by 21.3% in average precision.
Table 1: Zero-shot class-agnostic object detection performance of Grounding DINO [33] on MS-COCO [32], with hand-crafted prompts from various sources. We report average recall (AR) and precision (AP) limited to a maximum of 100 detections per image. $\Delta$AR quantifies the percentage decrease in AR comparing “query-merging” to “prediction-merging” for forming multi-word queries.
@@
Related Study. The full discussions can be found in Section A.1. Traditional bottom-up approaches for region proposal generation, such as those by [47] and [27], often face precision constraints despite high recall rates, limiting their scalability. Recent advancements in Vision Transformers (ViTs) by [4] and [10] have enabled self-supervised learning on massive datasets, extracting semantically meaningful features. Methods like LOST [45] and TokenCut [51] use graph-based techniques but are limited to detecting a single object per image. MOST [43] addresses this with entropy-based box analysis but struggles with generalization. MAVL [36] uses a late fusion strategy with text queries, requiring full supervision and multiple inferences. Our approach eliminates the need for labels and achieves state-of-the-art performance with one-pass inference using non-overlapping prompts. Vision-Language Models (VLMs), like those by [42] and [20], have shown potential in learning generic concepts. HierKD [35] and OV-DETR [58] align image representations with captions and extend DETR to open-vocabulary settings. GLIP [29], Grounding DINO [33], and T-Rex2 [21] integrate object detection and visual grounding. However, VLMs’ effectiveness depends on textual cues, and prompt tuning, as introduced by CoOp [24] and improved by CoCoOp and MaPLe [25], offers a solution by optimizing soft prompts while keeping the model’s parameters frozen. ProDA [34] learns diverse prompts using a Gaussian model. DFKD-VLFM [56] and PromptStyler [7] attempted to diversify a fixed number of prompts through contrastive approach. Despite these advancements, full supervision is typically required. UPL [17] and POUF [46] introduced unsupervised prompt learning, but adaptation for object detection remains limited. DiPEx is the first to apply prompt learning to class-agnostic object detection through a progressive self-training approach.
2 Pilot Study
In this section, we detail our preliminary exploration of the zero-shot detection capabilities using state-of-the-art VLM, Grounding DINO [33], to detect all objects irrespective of the associated classes on the MS-COCO dataset [32] as illustrated in Figure 1b. We conduct experiments using two types of text queries: Universal queries generated by ChatGPT for general object detection, and Class-Wide queries derived from WordNet, representing broad object categories. Our experiments reveal that semantic overlap between text queries impacts detection performance. To support this hypothesis, we conduct a case study showing that similar concatenated prompts reduce the model’s detection confidence.
2.1 Hand-crafted Queries for Class-agnostic Object Detection
@@
Universal Query. We employ ChatGPT to generate 13 synonyms of universal concepts, including nouns and adjectives, which are displayed as x-axis labels. The zero-shot object detection results, measured by average recall (AR) and precision (AP) across the top 100 confident boxes for each query text, are presented. The plot reveals that more general terms such as “generic” and “items” yield the highest AR. Surprisingly, more specific descriptors like “foreground”, “small”, or “tiny” tend to reduce AR and do not effectively aid in identifying foreground or small objects.
@@
Discussion on Multi-Word Queries. The zero-shot results presented in Figure 1b were obtained using single-word prompts for the Grounding DINO. To explore whether combining multiple words as prompts from a given source (e.g., WordNet) could improve zero-shot detection performance, we developed strategies for merging at both the input stage (query-merging) and the output stage (prediction-merging) as shown in Table 1. The query-merging strategy concatenates all input text queries (e.g., “foreground . elements . $\cdots$ tiny . objects .”) and performs a single-pass inference to obtain detections. The prediction-merging strategy, on the other hand, uses each text query individually for separate inference and then combines all box predictions. Table 1 shows that applying query-merging to Universal words results in a 52.46% reduction in AR compared to prediction-merging, whereas Class-wide queries (e.g., from WordNet) achieve a smaller decrease in AR of only 23.64%. These findings suggest that large semantic overlaps in concatenated queries (e.g., “stuff”, “objects” and “item” from ChatGPT) may greatly contribute to diminished object detection performance. To further investigate this phenomenon, we conducted a case study analyzing the impact of semantic overlap on detection performance, which is presented in the following section.
2.2 Confidence Diminishing when Text Query Semantically Overlap
To verify our hypothesis, we conduct a case study to demonstrate how semantic overlap in multi-word query leads to diminished detection confidence. We quantify semantic overlap by calculating the angular distance between pairs of textual token embeddings generated by BERT [9]. As shown in Figure 2, a small angular distance $\theta$ of 53.73 between the text tokens “plates” and “dishes” diminishes the model’s confidence. Consequently, some boxes that could be precisely localized with high confidence using the single token “plates” are omitted. In contrast, concatenating two text tokens with a larger angular distance (e.g., 60.99 between “plates” and “cup”) maintained high detection confidence. This combination resulted in bounding box predictions that encompassed all boxes predicted with each individual token (“plates” or “cup”). This case study supports our hypothesis that semantic overlap between concatenated text queries can interfere with the detection confidence of the model. Therefore, we propose that developing a method to learn a set of semantically non-overlapping prompts for the target dataset could enable efficient object localization with one-pass inference using VLMs.
@@ 3 Proposed Approach
In this section, we first mathematically formulate the task of class-agnostic detection using a general VLM and, without loss of generality, illustrate the process using Grounding DINO [33] as an exemplar model. We detail the steps of the proposed dispersing prompt expansion in Section 3.2, followed by the early termination strategy of the prompt set growth.
@@ 3.1 Problem Formulation
Class-agnostic OD. Let $\mathbf{I}$ denote the input image and $\mathbf{T}$ the associated text query. For the zero-shot object detection in a class-agnostic setting, we consider the text query $\mathbf{T}$ to be of the form of “a photo of a {class}”, where the class token {class} is sampled from our predefined Universal (e.g., “objects”) or Class-wide (e.g., “plant”) sets as described in Section 2. The text query is then tokenized and projected into word embeddings as $\mathbf{P}=\{\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{M},\mathbf{c}\}$, where $\mathbf{v}=\{\mathbf{v}_{i}\}_{i=1}^{M}\in\mathbb{R}^{M\times d}$ indicates a set of $M$ contextual embeddings and $\mathbf{c}$ is the query text embedding. Here, $d$ indicates the dimensions of learnable tokens. The visual embeddings $\mathbf{E}_{v}$ extracted from the visual encoder and prompt embeddings $\mathbf{P}$ are fused jointly to prompt the VLM and generate the final bounding box predictions $\mathbf{O}=f(\mathbf{E}_{v},\mathbf{P})\in\mathbb{R}^{N_{B}\times 4}$, with $f$ being the VLM, and $N_{B}$ being the number of predicted boxes. Formally, the objective of class-agnostic OD is to ensure that the generated bounding boxes can capture any objects as comprehensively as possible.